OpenStudy (anonymous):

What is the simplified form of 15 x to the fifth power over 24 y to the eighth power divided by 4 x squared over 8 y to the fourth power ? @midhun.madhu1987

OpenStudy (campbell_st):

well don't forget.... dividing by a fraction requires you to flip the 2nd fraction and multiply \[\frac{15x^5}{24y^8} \times \frac{8y^4}{4x^2} = \] use a system to simplify... numbers, letters, powers signs.. so you would have \[\frac{120x^4y^5}{96x^2y^8}\] now simplify by cancelling common factors